OpenStudy (anonymous):

Divide: 3x-21/x^2-49 divide 3x / x^2+7x

OpenStudy (anonymous):

\[\frac{3x-21 }{x^{2}-49 } \div \frac{3x }{x^{2}+7x }\]

OpenStudy (anonymous):

@Peter14 :)

OpenStudy (anonymous):

division of a fraction: invert and multiply then do it the same way as we did the problem earlier: factor everything, combine the fraction, then cancel.

OpenStudy (anonymous):

sorry, I have to go now.

OpenStudy (anonymous):

No problem, thanks!

hartnn (hartnn):

 \(\large \frac{3x-21 }{x^{2}-49 } \div \frac{3x }{x^{2}+7x }=..?\)